OpenStudy (anonymous):

A catapult launches a boulder with an upward velocity of 112 ft/s. The height of the boulder, h, in feet after t seconds is given by the function . How long does it take the boulder to reach its maximum height? What is the boulder’s maximum height? Round to the nearest hundredth, if necessary.

OpenStudy (anonymous):

is given by the 'what' function?

OpenStudy (anonymous):

oh sorry its y=-16t^2+112t+30

OpenStudy (cwrw238):

y = height?

OpenStudy (anonymous):

yes

OpenStudy (cwrw238):

differentiate to get expression for velocity ate time t at maximum height v = 0 so v = -32t + 112 = 0 solve for t

OpenStudy (cwrw238):

now plug in this value of t into your formula to get maximum height

OpenStudy (cwrw238):

do you follow that ok?

OpenStudy (cwrw238):

32t = 112 t = 3.5 secs = time to get maximum height maximum ht = -16(3.5)^2 + 112(3.5) + 30 ft

OpenStudy (cwrw238):

yes - same result as i got though they quoted the formula v = u + at

OpenStudy (anonymous):

u for initial velocity?

OpenStudy (cwrw238):

i derived v = u + at from your formula for the height at time t

OpenStudy (cwrw238):

u = 112 ft/s

OpenStudy (cwrw238):

yes u - initial velocity

OpenStudy (anonymous):

yeah guess they are pretty much the same, oh well :)

OpenStudy (cwrw238):

there's i believe 4 equations of motion for particle moving at constant velocity s = (u +v)t / 2 v = u + at s = ut + (1/2)at^2 v^2 = u^2 + 2as and are useful for problems of motion under gravity

OpenStudy (anonymous):

u is interchangeable with Vo where o is a subscript 0 what is s?

OpenStudy (cwrw238):

s is distance or more accurately displacement

OpenStudy (anonymous):

ahh okay, I know that as (triangle/delta) x or y depending on which vector, same thing though

OpenStudy (cwrw238):

e.g. if a projectile is dispatched from top of a building 100 ft high displacement is -100 - yea its a vector

OpenStudy (anonymous):